OpenStudy (anonymous):

Hi, could someone give me a step by step explanation of how to solve this problem? (3x)/(x+2) + (x)/(x-2) - (4)/(x²-4)

OpenStudy (anonymous):

So far this is what I have done:

OpenStudy (anonymous):

\[3x/(x+2) + x/(x-2) + -(4)/(x+2)(x-2)\] Made the common denominator- (x+2)(x-2) \[(3x)(x-2) + (x)(x+2) +(-4)(x-2)\] Multiply that out... \[3x²-6x+x²+2x+(-4x)-8\]

OpenStudy (anonymous):

Sorry, your question is unclear. What is the problem exactly?

OpenStudy (anonymous):

3x x 4x --- + --- - --- x+2 x-2 x²-4

OpenStudy (anonymous):

Those dashed lines are supposed to represent the lines in fractions. :/

OpenStudy (anonymous):

x²-4=(x+2)(x-2) You'll need to have the same denominator for all fractions to simplify the equation. 1=(x+2)/(x+2) eqn1 1=(x-2)(x-2) eqn2 3x/(x+2) =(3x/(x+2))*1 x/(x-2)=(x/(x-2))*1 4x/(x²-4) will remain the same since both (x+2) and (x-2) are factors of x²-4 Subsitute both 1s with eqn1 and eqn2 respectively to obtain the (x²-4) denominator. Simplify and you'll be able to get your answer. ;)

OpenStudy (anonymous):

Oh dear, the first time I did the problem I didn't multiply the 4x, but for some reason when I wrote it on here I had forgotten. Thank you for your help, but I have gotten to that point, and that is where the trouble starts. :/ I'll try working with it again, though.

OpenStudy (anonymous):

Sorry a little typing mistake: 1=(x-2)/(x-2) eqn2